A company wants to setup a new office in a country where the corporate tax rate is as follows: 15% of first $50,000 profits, 25% of next $25,000, 34% of next $25,000, and 39% of everything over $100,000. Executives estimate that they will have gross revenues of $500,000, total costs of $300,000, $30,000 in allowable tax deductions, and a one-time business start-up credit of $8000. What is taxable income for the first year, and how much should the company expect to pay in taxes?
1
Expert's answer
2015-10-12T09:20:24-0400
Taxable income for the first year will be: TI = 500,000 - 300,000 - 30,000 - 8,000 = $162,000. The company should expect to pay in taxes 15% of first $50,000 profits, 25% of next $25,000, 34% of next $25,000, and 39% of everything over $100,000, so it will pay T = 0.15*50,000 + 0.25*25,000 + 0.34*25,000 + 0.39*62,000 = $46,430.